OpenStudy (paul):

Solve for x. 4x + 6 = 38

OpenStudy (welshfella):

step 1 Subtract 6 from both sides of the equation step 2 Divide both sides by 4

OpenStudy (paul):

how?

OpenStudy (welshfella):

4x + 6 - 6 = 38 - 6 4x = 32 now dividing both sides by 4 will give you the value of x.

OpenStudy (paul):

ok

OpenStudy (welshfella):

so what do you think the value of x is?

OpenStudy (paul):

hold on

OpenStudy (paul):

i got 1 on 4/4 =1 and 32/4 =8

OpenStudy (paul):

is that right?

OpenStudy (paul):

what do you do with 1 and 8?

OpenStudy (paul):

i divided it and then what do you do with 1 and 8?

OpenStudy (photon336):

@paul can you show us your steps? your first objective should be to get x by itself

OpenStudy (welshfella):

4x / 4 = x 32/4 = 8

OpenStudy (welshfella):

@paul watch your language

OpenStudy (paul):

ya

OpenStudy (paul):

i know what is 32/4

OpenStudy (paul):

its 8

OpenStudy (paul):

and i wanna know what you have to do with 8

OpenStudy (welshfella):

right x = 8 is the solution

OpenStudy (paul):

thats it?

OpenStudy (welshfella):

yes solve for x means find the value of x

OpenStudy (paul):

ok thx

OpenStudy (welshfella):

yw

OpenStudy (paul):

what is yw?

OpenStudy (paul):

can you help me with one more?

OpenStudy (paul):

4 more

jhonyy9 (jhonyy9):

yw = you are welcome